OpenStudy (anonymous):

A parabola is defined by the following the following equation: x^(2)-10x-y+23=0 What are the coordinates of its vertex?

OpenStudy (amistre64):

gonna have to complete the square;

OpenStudy (amistre64):

x^2 -10x = y -23 x^2 -10x +25 = y-23 +25

OpenStudy (amistre64):

(x-5)^2 = (y+2)

OpenStudy (amistre64):

the vertex is at (5,-2) i believe

OpenStudy (amistre64):

teh equation tells us the vertex = (5,-2) the distance from vertex to focus or directix = 1/4 the focus is at (5,-1' 3/4) the parabola opens upwards

OpenStudy (anonymous):

everyone keeps telling me to complete the squares, but i don't even know how to do that. i am horrible at math.

OpenStudy (amistre64):

complete the square is a process used to make a square complete....add the missing parts really.

OpenStudy (amistre64):

A complete square can move freely back and forth between to forms; (x+3)^2 = x^2 +6x +9 ^^^^^^^^ this is a complete square

OpenStudy (amistre64):

if all we had was: x^2 +6x. then the square is missing a part to it and it is stuck in this form...

OpenStudy (amistre64):

How the incomplete square is made tells us alot about the missing parts

OpenStudy (amistre64):

in short, we need to take the number that is next to the 'x'.. that 6 there in this case; and split it in half

OpenStudy (amistre64):

6/2 = 3 the missing part that we need is then 3^2 x^2 +6x +3^2 = (x+3)^2

myininaya (myininaya):

THAN square so fun!

myininaya (myininaya):

well he already did it above so i was late

OpenStudy (amistre64):

lol..your just in time ;)

myininaya (myininaya):

:)

OpenStudy (amistre64):

chizz; it takes time and practice to know it good enough to be able to see it

myininaya (myininaya):

never talked about directex and focus in any of my math classes? what use are these?

OpenStudy (amistre64):

the focus and directix are geometry terms used to define a parabola; and any conic section really. It relies upon the geometric interpretation instead of the algebraic

myininaya (myininaya):

oh i see

OpenStudy (amistre64):

a parabola is: all poionts that are equi distant from a straight line from the focus and a perpendicular line to the directix.

OpenStudy (anonymous):

lol amistre, nice shirt? :D

OpenStudy (amistre64):

thnx :) i made it me self.

OpenStudy (anonymous):

seriously? ._. nice, lol np

OpenStudy (amistre64):

before i learned about openstudy, i taught myself many things; origami, knitting, crocheting, sister annoying...you know, the basics of life ;)

OpenStudy (anonymous):

sister annoying? that must've been an interesting experience lol. Your wife must be so grateful to have you >_<

OpenStudy (amistre64):

ironically no :/ but i did get to drive her to the hospital yesterday and see the kids for a bit :)

OpenStudy (anonymous):

nah, I bet she is :) but doesn't show it :P? that's nice ^_^